OpenStudy (anonymous):

A conveyor belt carries supplies from the first floor to the second floor, which is 12 feet higher. The belt makes a 60° angle with the ground. How far do the supplies travel from one end of the conveyor belt to the other? Round your answer to the nearest foot. (1 point) 21 ft 7 ft 14 ft 17 ft

OpenStudy (anonymous):

\[12 /x = \sin 60 \rightarrow x = 24/\sqrt{3} = 24 \sqrt{3}/ 3=8\sqrt{3}\] \[8\sqrt{3} \approx 13.8\] so i would say answer is 14 ft